# AED: Automatic Discovery of Effective and Diverse Vulnerabilities for Autonomous Driving Policy with Large Language Models

### v1

@@ Abstract
Assessing the safety of autonomous driving policy is of great importance, and reinforcement learning (RL) has emerged as a powerful method for discovering critical vulnerabilities in driving policies. However, existing RL-based approaches often struggle to identify vulnerabilities that are both effective—meaning the autonomous vehicle is genuinely responsible for the accidents—and diverse—meaning they span various failure types. To address these challenges, we propose AED, a framework that uses large language models (LLMs) to Automatically discover Effective and Diverse vulnerabilities in autonomous driving policies. We first utilize an LLM to automatically design reward functions for RL training. Then we let the LLM consider a diverse set of accident types and train adversarial policies for different accident types in parallel. Finally, we use preference-based learning to filter ineffective accidents and enhance the effectiveness of each vulnerability. Experiments across multiple simulated traffic scenarios and tested policies show that AED uncovers a broader range of vulnerabilities and achieves higher attack success rates compared with expert-designed rewards, thereby reducing the need for manual reward engineering and improving the diversity and effectiveness of vulnerability discovery.
I INTRODUCTION
Autonomous driving is a safety-critical domain where even minor decision-making flaws can lead to severe accidents [1]. One approach to assessing the safety of autonomous driving policies is vulnerability discovery, which systematically uncovers scenarios where autonomous driving policy fails or behaves undesirably [2, 3]. By controlling the behaviors of other traffic participants, vulnerability discovery create adversarial driving scenarios where the tested driving policy is exposed to hazardous conditions. Due to the rarity of such failure cases in real-world road tests, automated vulnerability discovery methods have been developed in simulation environments, where reinforcement learning (RL) emerges as an efficient approach that trains adversarial agents to induce accidents of the tested policy [4, 5, 6, 7].
@@ I INTRODUCTION
However, there are two key challenges in leveraging RL for vulnerability discovery. First, it is hard to automatically identify effective vulnerabilities that require determination of accident responsibility. When an accident occurs, it is essential to attribute the failure to the autonomous vehicle’s (AV) decision-making or those of other traffic participants. An effective vulnerability is one where the tested policy is responsible for the accident, revealing a genuine deficiency in its driving policy [8, 9, 7]. Accident responsibility, however, is complex to assess, often requiring human experts with domain knowledge in real-world settings [10, 11]. Similarly, in simulation-based RL training, designing a suitable reward function for responsibility attribution also demands substantial domain expertise [12, 13, 14], posing a challenge to the automated discovery of effective vulnerabilities.
The second challenge in RL-based vulnerability discovery is ensuring diverse failure cases. Since a safety-critical flaw can arise from various accident types, it is crucial to comprehensively explore the space of possible failures [9, 15, 7]. Traditional RL methods, however, tend to converge on the most probable failure case, leading to a lack of diversity in discovered vulnerabilities [16]. Existing approaches to diversity-aware RL generally fall into two types. The first type is optimizing a general diversity metric alongside task objectives. However, these methods often lead to vulnerabilities with uninterpretable differences that are not consistent with human understanding and regulations. The second type is incorporating heuristic-driven diversity measures tailored to specific tasks, which rely heavily on expert knowledge and struggle to generalize across different scenarios. These limitations hinder the interpretability and automation of diverse vulnerability discovery, leaving potential safety risks unsolved.
@@ II RELATED WORK
II-A Diverse Vulnerability Discovery
Adversarial RL has emerged as a popular approach to purposely generating safety-critical scenarios to efficiently find vulnerabilities where autonomous driving policies make mistakes. Kuutti et al. [4] use Advantage Actor Critic (A2C) [17] to induce rear-end collision. Some works [18, 5, 19] use Deep Q-Network [20] to generate corner cases to elicit AV mistakes. Chen et al. [6] use Deep Deterministic Policy Gradient (DDPG) [21] to create challenging lane-change conditions. Adaptive Stress Testing [9] incorporates a Responsibility-Sensitive Safety (RSS) [10] reward into Trust Region Policy Optimization (TRPO) [8] to classify AV failures with a single adversarial agent. Some works [22, 7] utilize Multi-agent DDPG [23] and Multi-agent Proximal Policy Optimization (PPO) [24] respectively to coordinate multiple adversarial agents. Some other studies [25, 26] use genetic algorithms to evolve adversarial agents to attack AV. However, these methods rely on manually designed reward functions, requiring extensive domain expertise to balance responsibility attribution and behavioral realism, which limits generalization across diverse traffic environments.
@@ II-A Diverse Vulnerability Discovery
To improve the diversity of discovered vulnerabilities, existing methods can be categorized into general and heuristic methods. General methods maximize information-theoretic objectives to encourage trajectory diversity [16] or measure geographic and semantic similarity between failure trajectories [27, 9]. However, these methods can lack interpretability for real-world accidents. Heuristic methods incorporate expert knowledge or discovered datasets to enhance diversity. Chen et al. [6] cluster multiple scenarios using DP-Means [28], while VDARS [7] proposes intrinsic rewards to promote novel failures. However, these methods still rely on manual design and domain-specific assumptions. Our work overcomes these limitations by using LLMs to automatically generate diverse reward functions for different vulnerability types, improving adaptability and reducing reliance on handcrafted heuristics.
II-B Reward Design with LLMs
Reward design is a fundamental challenge in reinforcement learning (RL), as it requires crafting reward functions that both reflect the intended behaviors and offer effective learning signals. Traditional reward engineering is often labor-intensive, relying heavily on domain expertise and iterative trial-and-error processes [29, 30]. With the rise of large language models (LLMs) pretrained on extensive human knowledge, their capabilities in reasoning, planning, and code generation have opened new possibilities for addressing core RL problems [31, 32, 33].
Among recent advances, utilizing LLMs for automatic reward modeling has become a promising direction. Some studies employ LLMs to interpret the environment and agent behavior, providing intrinsic rewards based on scenario understanding [12, 34, 35, 36]. Others focus on leveraging LLMs to generate executable reward code, often incorporating self-refinement mechanisms to iteratively improve the reward functions [13, 14, 37, 38]. Notably, Eureka [13] introduces an automatic reward design framework based on self-reflection. This approach leverages LLMs to generate executable reward functions using the environment’s source code and iteratively refines these functions based on evaluation results obtained from human-designed sparse rewards. Building on this concept, In-Context Preference Learning (ICPL) [14] extends the framework to tasks where human-designed rewards are unavailable, enabling LLMs to directly learn reward functions from human preferences. Our work extends ICPL to multi-agent and dynamic traffic scenarios, where reward design should account for complex agent interactions and responsibility attribution.
II-C Preference-Based Reinforcement Learning
@@
Preference-based reinforcement learning has been proposed to learn directly from an expert’s preference instead of handcrafted numeric reward functions, thereby reducing the reliance on extensive domain-specific prior knowledge. Studies [39, 40, 41, 42, 43, 44, 45] have explored RL from human feedback (RLHF) in the form of trajectory preference queries or score ranking to train a reward model. However, continuously querying human labels during online training is often impractical due to the associated cost and latency. To address this limitation, some studies have proposed RL from artificial intelligence feedback (RLAIF) [46]. Some works [47, 48] use Vision-Language-Models (VLMs) to align vision-based trajectory inputs with task-specific language descriptions. Some works [49, 50, 51] generate intrinsic reward functions using preference feedback from LLMs based on text descriptions of states or trajectories. Our work generates preference data pairs automatically using datasets produced by parallel adversarial policies. This enables scalable preference-based learning without numerous labelling while capturing diverse vulnerabilities under multi-agent interactions.
III PRELIMINARY
III-A Partially Observed Markov Decision Process
@@
Our MARL framework can be described as a Partially Observed Markov Decision Process (POMDP). defined by a tuple $\mathcal{M}=(\mathcal{N},\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{P},\mathcal{R},\gamma)$, where $\mathcal{N}=\{1,2,\cdots,N\}$ is the set of agents, $\mathcal{S}$ is the state space encompassing all possible states of the N agents, $\mathcal{A}$ is the shared action space, $\mathcal{O}$ denotes each agent’s observation space, $\mathcal{P}$ is the transition probability function, $\mathcal{R}$ is the joint reward function with $\gamma$ is the reward discount factor. At each step, agent $i$ receives an observation $o_{i}=\mathcal{O}(s_{i})$ where $s_{i}\in\mathcal{S}$ is the local state. The agent then takes an action $a_{i}$ according to the shared policy $\pi_{\theta}(a_{i}|o_{i})$, where $\theta$ represents the shared parameters. In the next step, the state progresses to $s_{i}^{\prime}$ according to the transition probability $p(s_{i}^{\prime}|s_{i},a_{i})\in\mathcal{P}$ and the agent $i$ receives a reward $r_{i}=\mathcal{R}(s_{i}|a_{i})$. For homogeneous agents, each agent $i$ optimizes the discounted cumulative rewards $J(\pi)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t}\gamma^{t}r_{t}\right]$, where a trajectory $\tau$ denotes a state-action-reward triplets sequence, i.e., $\tau=\{(s_{t},a_{t},r_{t})\}_{t}$.
III-B Effective and Diverse Vulnerability Discovery
@@
Our objective is to discover effective and diverse vulnerabilities in a fixed, pre-trained autonomous driving policy by training adversarial policies within the previously defined POMDP framework. While the tested driving policy remains unchanged, adversarial vehicle agents are trained to induce failure scenarios that expose weaknesses in the tested policy. Specifically, the adversarial policy is optimized to maximize the rates of effective vulnerability—defined as those for which responsibility can be attributed to the tested policy. The diversity is evaluated by the number of distinct vulnerability types, and the effectiveness is evaluated by the proportion of vulnerabilities discovered vulnerabilities attributable to the tested driving policy’s wrong decisions.
IV METHODOLOGY
@@
IV-A Automatic Reward Design
@@
To enable automatic reward design for vulnerability discovery, we extend the In-Context Preference Learning (ICPL) framework [14] to leverage the code-synthesis and in-context learning capabilities of LLMs. In our adaptation, the LLM generates task-specific reward functions based on environment descriptions and human preferences, guiding the discovery of effective accidents in a self-improving manner.
@@
The LLM first receives an environment description and a detailed task specification, then generates $K$ executable reward functions. The environment text outlines the scenario and relevant code-level details, such as the agent’s observation/action space and key parameters or functions useful for reward construction. The specific details of the task description are provided in Sec. IV-B. Each reward function is used to train an agent, resulting in $K$ agents, which are evaluated using human-written criteria for identifying effective accidents. To ensure comprehensive coverage, we also generate accident videos for human evaluation. Based on both metrics and visual inspection, evaluators select the best and worst reward functions. The selected best function typically yields high accident effectiveness across scenarios and serves as a basis for further refinement.
IV-B Diverse Accident Generation
@@
To enhance diversity in vulnerability discovery, the adversarial policy should capture distinct characteristics of various weaknesses in the tested policy. To achieve this, we introduce a Diverse Accident Generation component, which prompts LLMs to generate a set of reward functions that promote vulnerability diversity. By considering a wide range of accident types, LLMs automatically design multiple reward functions in parallel, each tailored to induce specific failure cases in the tested policy.
@@
The process begins with the LLM generating a task description for a given accident type based on the environment specifications. This description includes a high-level summary of the targeted accident scenario, concrete examples, justifications for attributing responsibility to the tested agent, and well-defined criteria for categorizing scenarios within this accident type. Using the component outlined in Sec. IV-A, the LLM then autonomously generates reward functions corresponding to different accident types in parallel.
@@
Once the reward functions are generated, they are integrated into the training of adversarial policies, each optimized for a specific vulnerability of the tested policy. These adversarial policies are then used to collect trajectory data, creating a structured dataset covering diverse failure cases. This dataset serves as a foundation for further analysis and improvement of the tested policy, ensuring a comprehensive evaluation of its robustness across various accident scenarios.
IV-C Preference-Based Effectiveness Enhancement
@@
Reward functions derived from Sec. IV-B for a specific accident type are often imperfect, sometimes capturing unintended accident types or ineffective scenarios. For instance, the reward function for Left Lane Change (Accident Type 1) may capture Right Lane Change (Accident Type 2) and Adversarial Vehicle cut in mistakenly (Accident Type 3). To improve the effectiveness and distinguish diverse accident types, we propose a preference-based effectiveness enhancement component to refine the LLMs-driven reward function. The key idea is to learn a new reward function $\hat{r}$ that prefers trajectories corresponding to the intended accident types over other types. Specifically, $\hat{r}$ is trained to assign higher cumulative rewards to Type 1 than to those from Type 2 or Type 3.
@@
Figure 3: Number of vulnerability types discovered across different environments, tested policies, and numbers of adversarial vehicle agents. Each line represents the cumulative number over three training iterations, where the first discovered type is counted as 1, and subsequent new types are weighted by their relevant frequency among effective failures. AED consistently discovers more diverse vulnerabilities than VDARS.
@@
In this way, the reward function $\hat{r}$ learns to favor Type 1 accident trajectories while effectively distinguishing them from other types. We then incorporate $\hat{r}$ and the original LLMs-driven reward function into RL training. If $\hat{r}$ identifies trajectories as Type 2, the training will be pernalized. This guides RL training to prioritize Accident Type 1 and ultimately improve the attack success rate for this category.
V EXPERIMENT
We evaluate AED in two driving environments—Highway and Roundabout—using the Highway simulator [52]. We consider three types of tested driving policy: two Planning-based methods, Value Iteration (VI) [53] and Robust Value Iteration (RVI) [54], and one RL-based method, Dueling Double Q-Network (D3QN) [55]. To account for the complexity of traffic interactions, we conduct experiments with two and three adversarial vehicle agents in each environment.
V-A Implementation Details
We consider three distinct types of vulnerabilities in our experiments: Left Lane Change (Left), Right Lane Change (Right), and Rear End (Rear) collision scenarios. These categories correspond to the different failure-inducing situations: an accident caused during a left lane change maneuver, a right lane change maneuver, or a rear-end collision. To ensure robustness, each adversarial training is repeated with three different random seeds. Each adversarial policy is trained using the MAPPO algorithm [24] for 10 million steps. We utilize GPT-4o for diverse accident generation.
@@ V-A Implementation Details
  • VDARS [7]: A multi-agent RL framework that employs handcrafted reward functions to learn the adversarial policy for vulnerability discovery.

### v2

@@ Abstract
Assessing the safety of autonomous driving policy is of great importance, and reinforcement learning (RL) has emerged as a powerful method for discovering critical vulnerabilities in driving policies. However, existing RL-based approaches often struggle to identify vulnerabilities that are both effective—meaning the autonomous vehicle is genuinely responsible for the accidents—and diverse—meaning they span various failure types. To address these challenges, we propose AED, a framework that uses large language models (LLMs) to Automatically discover Effective and Diverse vulnerabilities in autonomous driving policies. We first utilize an LLM to automatically design reward functions for RL training. Then we let the LLM consider a diverse set of accident types and train adversarial policies for different accident types in parallel. Finally, we use preference-based learning to filter ineffective accidents and enhance the effectiveness of each vulnerability. Experiments across multiple simulated traffic scenarios and tested policies show that AED uncovers a broader range of vulnerabilities and achieves higher attack success rates compared with expert-designed rewards, thereby reducing the need for manual reward engineering and improving the diversity and effectiveness of vulnerability discovery. The implementation can be found on: https://github.com/thu-nics/AED.
INTRODUCTION
Autonomous driving is a safety-critical domain where even minor decision-making flaws can lead to severe accidents (guo2019safe). One approach to assessing the safety of autonomous driving policies is vulnerability discovery, which systematically uncovers scenarios where autonomous driving policy fails or behaves undesirably (rastogi2020threats; wan2022too). By controlling the behaviors of other traffic participants, vulnerability discovery create adversarial driving scenarios where the tested driving policy is exposed to hazardous conditions. Due to the rarity of such failure cases in real-world road tests, automated vulnerability discovery methods have been developed in simulation environments, where reinforcement learning (RL) emerges as an efficient approach that trains adversarial agents to induce accidents of the tested policy (kuutti2020training; feng2021intelligent; chen2021adversarial; mu2024multi).
@@ INTRODUCTION
However, there are two key challenges in leveraging RL for vulnerability discovery. First, it is hard to automatically identify effective vulnerabilities that require determining accident responsibility. When an accident occurs, it is essential to attribute the failure to the autonomous vehicle’s (AV) decision-making or those of other traffic participants. An effective vulnerability is one where the tested policy is responsible for the accident, revealing a genuine deficiency in its driving policy (schulman2015trust; corso2019adaptive; mu2024multi). Accident responsibility, however, is complex to assess, often requiring human experts with domain knowledge in real-world settings (shalev2017formal; sun2024acav). Similarly, in simulation-based RL training, designing a suitable reward function for responsibility attribution also demands substantial domain expertise (kwon2023reward; ma2023eureka; yu2025largelanguagemodelsincontext), posing a challenge to the automated discovery of effective vulnerabilities.
The second challenge in RL-based vulnerability discovery is ensuring diverse failure cases. Since a safety-critical flaw can arise from various accident types, it is crucial to comprehensively explore the space of possible failures (corso2019adaptive; koren2020adaptive; mu2024multi). Traditional RL methods, however, tend to converge on the most probable failure case, leading to a lack of diversity in discovered vulnerabilities (eysenbach2018diversity). Existing approaches to diversity-aware RL generally fall into two types. The first type is optimizing a general diversity metric alongside task objectives. However, these methods often lead to vulnerabilities with uninterpretable differences that are not consistent with human understanding and regulations. The second type is incorporating heuristic-driven diversity measures tailored to specific tasks, which rely heavily on expert knowledge and struggle to generalize across different scenarios. These limitations hinder the interpretability and automation of diverse vulnerability discovery, leaving potential safety risks unsolved.
@@
Diverse Vulnerability Discovery
Adversarial RL has been widely used to generate safety-critical scenarios that reveal vulnerabilities in autonomous driving policies. Prior works use Deep Q-Network (mnih2015human) to generate corner cases (karunakaran2020efficient; feng2021intelligent; sun2021corner), and apply DDPG (lillicrap2015continuous) to create challenging lane-change conditions(chen2021adversarial). Adaptive Stress Testing (corso2019adaptive) incorporates a Responsibility-Sensitive Safety (RSS) (shalev2017formal) reward into TRPO (schulman2015trust) to classify AV failures. Multi-agent extensions such as MADDPG (lowe2017multi) and MAPPO (yu2022surprising) have been applied to coordinate multiple adversarial vehicles (wachi2019failure; mu2024multi). However, these methods rely on crafted reward functions, which require domain expertise to balance responsibility attribution and behavioral realism.
@@
To improve the diversity of discovered vulnerabilities, existing methods can be categorized into general and heuristic methods. General methods optimize information-theoretic objectives (eysenbach2018diversity) or apply trajectory similarity metrics (liu2012similarity; corso2019adaptive), but often lack interpretability in real-world scenarios. Heuristic methods incorporate expert knowledge, such as clustering (chen2021adversarial) or intrinsic novelty reward (mu2024multi) to promote rare failures, but remain dependent on handcraft designs and domain-specific assumptions. Our work addresses these limitations by using LLMs to automatically generate diverse reward functions tailored to distinct vulnerability types, improving adaptability while reducing the reliance on manual heuristics.
Reward Design with LLMs
Reward design is a fundamental challenge in reinforcement learning, often requiring extensive domain expertise and iterative trial-and-error (sutton2018reinforcement; singh2009rewards). Recent advances leverage LLMs to automate reward design by exploiting their capabilities in reasoning, planning, and code generation (shinn2023reflexionlanguageagentsverbal; du2023guidingpretrainingreinforcementlearning; carta2024groundinglargelanguagemodels). Some studies use LLMs to interpret environment and agent behaviors and assign intrinsic rewards based on scenario understanding (kwon2023reward; wu2024readreaprewardslearning; chu2023accelerating; wang2024sociallymorallyawarerl), while others focus on generating executable reward code (ma2023eureka; yu2025largelanguagemodelsincontext; yu2023languagerewardsroboticskill; li2024automcrewardautomateddense). Notably, Eureka (ma2023eureka) proposes an automatic self-reflective framework where LLMs generate and iteratively refine reward functions using environment code and sparse human-designed evaluation rewards. ICPL (yu2025largelanguagemodelsincontext) extends this to tasks without predefined rewards, enabling preference-based learning from human feedback. Our work extends ICPL to multi-agent traffic scenarios, where reward design must consider agent interactions and responsibility attribution.
Preference-Based Reinforcement Learning
@@
Preference-based RL learns directly from expert preference rather than handcrafted numeric rewards, reducing the reliance on domain-specific knowledge. Prior works (akrour2012april; wilson2012bayesian; wirth2016model; mathewson2017actor; lee2021pebble; myers2022learning; christiano2017deep) explore RL from human feedback via trajectory preference queries or score ranking. However, querying human labels during online training is often costly and slow. To address this limitation, recent studies have proposed RL from AI feedback (bai2022constitutional). Some leverage LLMs to provide preference based on vision-based trajectory or text descriptions of states and actions (wang2024rl; wang2025prefclm; klissarov2023motif; chu2023accelerating; tu2024online). Our work automatically generates preference data pairs from simulation outcomes produced by parallel adversarial policies, enabling scalable preference learning without numerous labeling while capturing diverse vulnerabilities in multi-agent interactions.
PRELIMINARY
Partially Observed Markov Decision Process
@@
Our MARL framework is modeled as a Partially Observed Markov Decision Process (POMDP). defined by a tuple $\mathcal{M}=(\mathcal{N},\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{P},\mathcal{R},\gamma)$, where $\mathcal{N}=\{1,2,\cdots,N\}$ is the set of agents, $\mathcal{S}$ is the state space covering all possible states of the N agents, $\mathcal{A}$ is the shared action space, $\mathcal{O}$ denotes each agent’s observation space, $\mathcal{P}$ is the transition probability function, $\mathcal{R}$ is the joint reward function with $\gamma$ is the reward discount factor. At each step, agent $i$ receives an observation $o_{i}=\mathcal{O}(s_{i})$ where $s_{i}\in\mathcal{S}$ is the local state. The agent then takes an action $a_{i}$ according to the shared policy $\pi_{\theta}(a_{i}|o_{i})$, where $\theta$ represents the shared parameters. The next state progresses to $s_{i}^{\prime}$ according to the transition probability $p(s_{i}^{\prime}|s_{i},a_{i})\in\mathcal{P}$ and the agent $i$ receives a reward $r_{i}=\mathcal{R}(s_{i},a_{i})$. For homogeneous agents, each agent $i$ optimizes the discounted cumulative rewards $J(\pi)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t}\gamma^{t}r_{t}\right]$, where a trajectory $\tau$ denotes a state-action-reward triplets sequence, i.e., $\tau=\{(s_{t},a_{t},r_{t})\}_{t}$.
Diverse and Effective Vulnerability Discovery
@@
Our objective is to discover diverse and effective vulnerabilities in a fixed, pre-trained autonomous driving policy by training adversarial policies within the previously defined POMDP framework. While the tested policy remains fixed, adversarial vehicles are optimized to induce failure scenarios that expose weaknesses in the tested policy. Specifically, the adversarial policy is optimized to maximize the rate of effective vulnerability, defined as those for which responsibility can be attributed to the tested policy. Diversity is assessed by the number of distinct vulnerability types, and effectiveness is assessed by the proportion of vulnerabilities discovered arising from wrong decisions of the tested driving policy. Formal definitions and evaluation procedures for these two metrics are provided in Sec. Evaluation of Diversity and Sec. Evaluation of Effectiveness.
METHODOLOGY
@@
Automatic Reward Design
@@
To enable automatic reward design for vulnerability discovery, we extend the In-Context Preference Learning (ICPL) framework (yu2025largelanguagemodelsincontext) to leverage the code-synthesis and in-context learning capabilities of LLMs. In our adaptation, the LLM generates task-specific reward functions based on environment descriptions and human preferences, guiding the discovery of effective accidents in a self-improving manner.
@@
The LLM first receives an environment description and a detailed task specification, then generates $K$ executable reward functions. The environment text outlines the scenario and relevant code-level details, such as the agent’s observation/action space and key parameters or functions useful for reward construction. The specific details of the task description are provided in Sec. Diverse Accident Generation. Each reward function is used to train an agent, resulting in $K$ agents, which are evaluated using human-written criteria for identifying effective accidents. To ensure comprehensive coverage, we also generate accident videos for human evaluation. Based on both metrics and visual inspection, evaluators select the best reward functions. The selected best function typically yields high accident effectiveness across scenarios and serves as a basis for further refinement.
Diverse Accident Generation
@@
To enhance diversity in vulnerability discovery, the adversarial policy should capture distinct characteristics of various weaknesses in the tested policy. We introduce a Diverse Accident Generation component, which prompts LLMs to generate a set of reward functions that promote vulnerability diversity. By considering a wide range of accident types, LLMs automatically design multiple reward functions in parallel, each tailored to induce specific failure cases.
@@
The process begins with the LLM generating a task description for a given accident type based on environment specifications. This description includes a high-level summary of the targeted accident scenario, concrete examples, justifications for attributing responsibility to the tested agent, and criteria for categorizing scenarios within this accident type. Using the component outlined in Sec. Automatic Reward Design, the LLM then autonomously generates reward functions corresponding to different accident types in parallel.
@@
Once reward functions are generated, they are integrated into the training of adversarial policies, each optimized for a specific vulnerability of the tested policy. These adversarial policies are then used to collect trajectory data, creating a structured dataset covering diverse failure cases. This dataset serves as a foundation for further analysis and improvement of the tested policy, ensuring a comprehensive evaluation of its robustness across various accident scenarios.
The prompts used to guide the LLM in generating diverse accident types, along with the corresponding environment and scenario descriptions, are provided in Appendices.
Figure 3: Demonstration of “Highway” and “Roundabout”.
Figure 4: Examples of distinct vulnerability types discovered by AED. Adv. is abbreviation for “adversarial”.
Figure 5: Number of vulnerability types discovered across different environments, tested policies, and numbers of adversarial vehicle agents. Each line represents the cumulative number over six training iterations, where the first discovered type is counted as 1, and subsequent new types are weighted by their relevant frequency among effective failures. AED consistently discovers more diverse vulnerabilities than VDARS.
Preference-Based Effectiveness Enhancement
@@
Reward functions derived from Sec. Diverse Accident Generation for a specific accident type are often imperfect, sometimes capturing unintended accident types or ineffective scenarios. For instance, the reward function for Left Lane Change (Accident Type 1) may capture Right Lane Change (Accident Type 2) and Adversarial Vehicle cut in mistakenly (Accident Type 3). To improve the effectiveness and distinguish diverse accident types, we propose a preference-based effectiveness enhancement component to refine the LLMs-driven reward function. The key idea is to learn a new reward function $\hat{r}$ that prefers trajectories corresponding to the intended accident types over other types. Specifically, $\hat{r}$ is trained to assign higher cumulative rewards to Type 1 than to those from Type 2 or Type 3.
@@
In this way, the reward function $\hat{r}$ learns to favor Type 1 accident trajectories while effectively distinguishing them from other types. We then incorporate $\hat{r}$ and the original LLMs-driven reward function into RL training. If $\hat{r}$ identifies trajectories as Type 2, the training will be penalized. This guides RL training to prioritize Accident Type 1 and ultimately improve the attack success rate for this category.
EXPERIMENT
We evaluate AED in two driving environments “Highway” and “Roundabout” using the Highway simulator (leurent2018environment), as shown in Fig. 3. We consider three types of tested driving policy: two Planning-based methods, Value Iteration (VI) (bellman1966dynamic) and Robust Value Iteration (RVI) (nilim2004robust), and one RL-based method, Dueling Double Q-Network (D3QN) (wang2016dueling). To account for the complexity of traffic interactions, we conduct experiments with two and three adversarial vehicles in each environment.
Implementation Details
We categorize vulnerabilities into three types based on the relative positions of the tested and adversarial vehicles: Left Lane Change (Left), Right Lane Change (Right), and Rear End (Rear) collision scenarios. These categories correspond to different failure-inducing situations. For later diversity analysis, we further divide these categories into finer-grained subtypes based on common real-world traffic accident patterns. Each adversarial policy is trained using the MAPPO algorithm (yu2022surprising) for 10 million steps. We utilize GPT-4o and DeepSeek-R1 for diverse accident generation. Detailed environment settings and training parameters are provided in the Appendices.
@@
  • VDARS (mu2024multi): A multi-agent RL framework that employs handcrafted reward functions to learn the adversarial policy for vulnerability discovery.